OpenStudy (anonymous):

10). Multiply the following expressions. Be sure to reduce all answers to lowest terms. (The numerator and denominator of the answer should not have any factors in common.) d + y / 3 multiply by 12 / d + y

OpenStudy (anonymous):

i think you need some parenthesis. ((d+y)/3)*(12/(d+y)). d+y cancels, and we get 12/3=(4*3)/(1*3)=4/1=4